# Extracting Affect Aggregates from Longitudinal Social Media Data with Temporal Adapters for Large Language Models

### v1

@@ Introduction
Temporal Adapter Fine-Tuning
@@
We select the base pretrained version (i.e., not instruction tuned) of a high-performing and openly available model, Llama 3 8B (AI@Meta 2024), to fine-tune the model on plain text data. We fine-tune Temporal Adapters with the causal language modeling objective, i.e. predicting next tokens, as shown in Figure 2. Adapters are a parameter-efficient approach to fine-tuning LLMs that allows researchers to easily swap model properties (Pfeiffer et al. 2023). LoRA adapters (Hu et al. 2021) add rank-decomposition matrices to each layer of the transformer architecture. When fine-tuning an LLM with LoRA adapters, all original model weights are frozen and only the added weights are trained. We train LoRA adapters of rank $128$ as a compromise between fine-tuning efficiency and training quality. This is twice the amount of parameters as the authors of the original LoRA paper considered necessary (Hu et al. 2021), but still only $0.67\%$ of the weights that would be trained when fully fine-tuning the model.
@@
We conduct a partial hyperparameter search by training adapters of varying rank, with different learning rates, and for up to 8 epochs. We find that fine-tuning adapters works best for our purposes with a small learning rate of $5*10^{-6}$ and at most $1$ training epoch. The training loss is mostly stable after $0.5$ epochs at an average of $3.4$ across all adapters. In a preliminary experiment with a larger learning rate of $10^{-4}$, training and test set loss decreased to around $2.6$ after $8$ training epochs but the cross-correlations of the attitudes we extracted with survey data were much lower. This is most likely due to the model over-fitting on the training data (tweets) that is quite different from the survey questions we prompt at inference time. We train adapters with 3 different training seeds for each week to better understand the reliability of our training method. We used bfloat16 and the adamw_torch_fused optimizer with a batch size of $6$ and $4$ gradient accumulation steps. The training was conducted on two NVIDIA H100 GPUs in Distributed-Data-Parallel mode and takes approximately 20 minutes per adapter.
Extracting Survey Answers
@@
Once we have finished adapter training, we focus on model inference and extract answers to survey questions as follows. We first concatenate a survey question with each of its $n$ answer options, obtaining $n$ separate prompts. We experiment with an optional answer prefix that is build for the particular survey question. For instance, for Britain’s Mood, Measured Weekly, we add the optional answer prefix ‘I felt’ which is followed by the answer options ‘happy’, ‘sad’ etc. This is to accommodate for the fact that first-token probabilities can be unreliably for scoring survey answers from LLMs (Wang et al. 2024b). We prompt each answer option separately and without additional labels as previous research has shown order effects and the tendency for LLMs to prefer certain answer labels (Tjuatja et al. 2024; Dominguez-Olmedo, Hardt, and Mendler-Dünner 2023; Wang et al. 2024b).
@@
Based on a survey answer scoring method used in previous research (Wu et al. 2024; Naous et al. 2024), we perform a single forward pass of the concatenated question, optional answer prefix, and answer option through the LLM. We gather the token probabilities from the last LLM layer for each token in the answer option after applying softmax. If an answer option consists of multiple tokens, we multiply their probabilities. We further vary temperature, a normalization parameter applied to the final softmax function, between $0.25$ and $4$. Our extraction method is deterministic, i.e., if the weights of an LLM are kept constant, so will be the answer probabilities we extract. Unless otherwise noted, we report results for temperature $1$, i.e., standard softmax.
@@
We perform survey answer extraction with each weekly adapter separately activated in the LLM and for each of the $3$ seeds. Since swapping adapters and inference is computationally inexpensive, we extract survey answers after training each adapter for $1$ epoch, as well as after every $50$ steps of training. We obtain a time series for each combination of question, answer options, and set of inference hyperparameters. YouGov presents weekly estimates using a smoothed trend line to reduce random fluctuations due to sampling variability. We follow this approach and apply a $3$ week rolling average. For plotting time series, we apply min-max normalization and show the mean probability for each extracted survey answer across all $3$ seeds.
Comparison and Validation
@@
Figure 3: Affect Aggregates Extracted from Temporal Adapters. We extract answer probabilities by prompting a weekly fine-tuned Llama 3 8B with the same question wording as in the survey (YouGov 2024a), and compare them to the respective weekly survey data. The time series are min-max normalized and a $3$ week rolling average is applied. The shaded orange area indicates minimum and maximum LLM answer probabilities across $3$ training seeds. Our results descriptively show in the plot a similar trend of both signals and we find strong positive and significant ($p<0.01$) cross-correlation between LLM probabilites and the survey data. Additional time series are provided in Figures 7 and 8 in the Appendix.
@@
We evaluate our longitudinal macroscopes of mood and attitudes by cross-correlating them with the respective survey data using Pearson’s correlation coefficient. YouGov publishes weekly survey results on Britain’s Mood, which results in $35$ data points in our observation period that we can use to compute cross-correlation. Results on British attitude towards the NHS are only published monthly, i.e., we are left with 9 relevant data points. We perform permutation tests with $10.000$ permutations of our time series, while keeping the survey data intact, to obtain a significance value for our cross-correlations.
@@
We compare our results to emotion aggregates from a BERT-based model for emotion detection, TweetNLP (Camacho-Collados et al. 2022). This baseline model will help us differentiate between issues in our training data, e.g. coverage error, or differences in expression of emotion, and measurement error specific to our research setup. TeetNLP was pre-trained on the tweet_eval dataset that consists of tweets labeled with $11$ emotions (Mohammad et al. 2018). The model achieves a macro F1 score of $0.72$ on the tweet_eval test set. We classify each tweet in our weekly training datasets separately and calculate the share of every emotion in every week. Being pre-trained on Twitter data makes TweetNLP particularly suited as a baseline model for our setup. However, only 4 of the emotions classified by TweetNLP match with emotions included in Britain’s Mood, Measured Weekly: ‘fear’/‘scared’, ‘sadness’/‘sad’, ‘joy’/‘happy’, and ‘optimism’/‘optimistic’. We again cross-correlate the results obtained from TweetNLP with the survey data and perform a permutation test.
Synthetically Mixed Data
To explore the internal validity of our mood macroscopes, we perform additional experiments with synthetically mixed data that is labeled as ‘happy’ or ‘sad’. With this setup, we can investigate whether a different prevalence of emotion signals indeed corresponds to different affect aggregates obtained from our method. We hypothesize a linear relationship between the amount of ‘sad’ tweets in the training data and the token probability for ‘sad’ as an answer option, and vice versa for ‘happy’. There are $1163$ tweets labeled ‘happy’ and $1326$ tweets labeled ‘sad’ in the tweet_eval dataset. From these tweets, we select $1163$ using $11$ splits: $100\%$ happy + $0\%$ sad, $90\%$ happy + $10\%$ sad, …, and $0\%$ happy + $100\%$ sad. We then train adapters on each of these splits using the same hyperparameters as described above. We repeat the splitting and training procedure using $10$ random seeds to improve robustness. Since the synthetically mixed training datasets are much smaller than our weekly Twitter datasets, the number of training steps/epochs is not directly comparable.
Results
@@
Figure 3 shows time series for two collective emotions, scared and happy, that we extract using weekly Temporal Adapters. We use the original survey question that was developed by YouGov (2024a) and compare our results to nationally representative survey data. We cap the training steps such that we have the same amount of training data across weeks. The plot shows the mean answer probability across $3$ training seeds, with minimum and maximum probabilities, after a $3$ week rolling average was applied and after the time series were min-max normalized. Time series for other emotions are provided in Figures 7 and 8 in the Appendix.
@@
Overall, we observe similar trends between our estimates of collective emotion and the survey data. Scared spikes on the March 23rd, 2020, when the first UK lockdown in response to the COVID-19 pandemic was announced [^2] https://www.theguardian.com/world/2020/mar/23/boris-johnson-orders-uk-lockdown-to-be-enforced-by-police, Accessed Sept 8th, 2024. Another increase can be seen in our estimate of collective fear around June 1st 2020, at a time when lockdown restrictions were partially lifted333for England, see for instance https://www.legislation.gov.uk/uksi/2020/558/made, Accessed Sept 8th, 2024. Fear also sees an increase around Christmas 2019. The survey data similarly indicates the largest level of collective fear around the announcement of the lockdown and a small increase around Christmas, but the second spike around June 1st, 2020 is not observed in the survey. This is likely due to sampling errors in the social media data, i.e., more concerned people being active on Twitter, or due to differences in the expression of emotions on social media and in surveys. TweetNLP shows the same second spike in 2020 when used to classify the same training data, as shown in Appendix Figure 8.
For collective happiness, we similarly extract estimates with good face validity. Happiness decreases rapidly in the two weeks before the nationwide lockdown was announced and recovers after another two weeks, presumably when people acclimatized with the new situation. This follows the trend that we observe in survey data. Again, our estimate deviates from survey data around June 1st 2020, when restrictions were partially lifted.
@@
Compared to TweetNLP, which extracts the lowest happiness after June 1st, however, our method’s estimate is closer to the YouGov survey data in that it identifies the lowest happiness around March 23rd. It should be addressed that we observe larger confidence intervals across $3$ training seeds, and larger week-by-week fluctuation for ‘happy’ as compared to ‘scared’. One possible explanation for this would be that there is less information about happiness in our training data, which would lead to less numerical influence in adapter fine-tuning, and thus result in larger confidence intervals after min-max scaling. Another explanation could be that information about happiness is less evenly distributed in our training data. We sample an equal amount of text for each week and randomly shuffle the training data, so less even distribution of information about happiness would increase the relevance of different training seeds. Thirdly, the LLM could be less capable of learning information about happiness during training, leading to more error surrounding the concept.
Taking a step back, we calculate cross-correlations (Pearson) between the longitudinal data we extract and the respective survey data, and perform a significance test with $10,000$ permutations. Figure 4 shows the results for all emotions found in Britain’s Mood, Measured Weekly across 3 training seeds, as well as the cross-correlation for our baseline method, TweetNLP (Camacho-Collados et al. 2022). We observe three groups of emotions: (i) emotions that show a strong positive, significant correlation, (ii) emotions that are weakly positively or negatively correlated, and (iii) optimism, which is strongly negatively correlated. In the first group we find both positively and negatively connoted emotions, and in particular affective phenomena that are sometimes referred to as being part of a set of basic emotions: ‘happiness’, ‘fear’, and ‘sadness’ (Ekman 1992). For collective emotions where an estimate can be extracted from TweetNLP, i.e., where labeled training data was available, we find that our method produces comparable results. We notably also obtain strong positive, significant correlations for emotions for which no TweetNLP estimate is available: ‘frustration’, ‘boredom’, and ‘being energetic’. For group (ii), is remains unclear why these collective emotions are hard to estimate from the data that we have available. One possible explanation might be that in contrast to the affective phenomena found in group (i), group (ii) contains more phenomena that clearly fall under ‘mood’ rather than ‘emotion’, i.e., they are more diffuse and longer lived, which might contradict typical social media behavior. Finally, we find a significant strong negative cross-correlation for ‘optimistic’, both for our estimate as well as for the baseline model. This points to general differences in the expression of this emotions between surveys and social media.
@@
Figure 4: Several Extracted Emotions Highly Correlate with Survey Data. We cross-correlate the answers we extract from Llama 3 with the respective British survey data (YouGov 2024a). Across $3$ training seeds, we show minimum and maximum correlation with error bars and indicate the worst $p$ value (*$p<0.05$,**$p<0.01$,***$p<0.001$). Our results vary strongly between emotions. They are in line with the baseline model’s estimates, and extend them to additional emotions by not requiring labeled training data.
@@
We create synthetically mixed training data from the tweet_eval dataset (Mohammad et al. 2018) in 11 splits ranging from $100\%$ sad to $100\%$ happy. We then prompt the model with the same question as used in Britain’s Mood, Measured Weekly and extract answer probabilities for the answer options ‘happy’ and ‘sad’. We hypothesize a positive linear correlation between the share of ‘happy’ tweets in the training data and the probability for answer option ‘happy’, and vice versa for ‘sad’. Figure 5 shows the mean and standard deviation of extracted answers across 10 training seeds. We find strong positive correlations for both answer options, which supports the internal validity of our method – more ‘happy’ tweets actually lead to a higher answer probability for ‘happy’ and vice versa for ‘sad’. For both emotions, the relationship between training data and extracted answer is surprisingly linear. Our answer scoring method does not ensure that the probabilities across all answer options add up to $1$, since we apply softmax on the last layer of the LLM, and each answer probability depends on probabilities of all other tokens in the vocabulary. The wide error bands indicate large random error due in training and underline the importance of evaluating results obtained from multiple seeds.
@@
Figure 5: Internal Validity Demonstrated in Experiments with Synthetically Mixed Data. We synthetically mix LLM training data with splits ranging from data that is labeled $100\%$ sad to $100\%$ happy. We then extract answers to YouGov (2024a)’s survey question at each split, and show mean and standard deviation over 10 training seeds. The results support the internal validity of our extraction method, but also highlight random error in training and a nonlinear relationship between training data ratio and extracted estimate.
(a) Robustness Across Multiple Survey Instruments: YouGov & PANAS-X
(b) Application to Attitude Aggregates
Figure 6: We extract the same collective emotions with (6(a)) an additional survey instrument (PANAS-X, Clark and Watson 1994) and (6(b)) separately extract a collective attitude towards the NHS. We cross-correlate our results with the respective YouGov survey data. We find that our method is robust across multiple survey instruments and can be applied to the extraction of collective attitudes as well. Across $3$ training seeds, we indicate minimum and maximum correlation with error bars and show the worst $p$ value (*$p<0.05$,**$p<0.01$,***$p<0.001$). For time series, see Figures 9 and 10 in the Appendix.
@@
In preliminary experiments, we investigated different hyperparameters for model training and for the extraction of affect aggregates, see Figure 11 in the Appendix. We investigated cross-correlations after different numbers of training steps and epochs and found little benefit in fine-tuning the model for more than $1$ epoch. While longer fine-tuning further improves training loss in the causal language modeling objective, the highest cross-correlations with survey data are observed relatively early on. This is likely due to the language model overfitting on the training data, which itself is quite different from the questions that we use for answer extraction. Similar also applies to learning rate – we experimented with learning rates between $10^{-3}$ and $10^{-6}$ and found that smaller learning rates generally perform better. Only $10^{-6}$ was too small, as the training loss remained almost constant, so we opted for $5*10^{-6}$ instead. Finally, we found that training all adapters with the same amount of training data generally works better than training all adapters for $1$ epoch on all the data that was available in this week. In other words, we extracted answers for all Figures presented previously in this section after training each adapter for $350$ steps instead of $1$ epoch. For answer extraction, we investigated the effect of temperature and of whether or not an answer prefix was attached to the answer options. A lower temperature leads to less evenly distributed token probabilities, similar to ‘enhancing contrast’. We found that this is beneficial when extracted attitudes have low noise, but that it can also increase random fluctuation. We also found that lower temperatures tend to overemphasize some answer options. Finally, adding an answer prefix lead to more consistent results. This is in line with previous findings on first-token probabilities (Wang et al. 2024b) and is likely related to LLMs being trained to create text rather than to answer surveys with only an answer option.

### v2

@@ 4 Temporal Adapters for LLMs
4.1 Temporal Adapter Fine-Tuning
@@
We select the base pretrained version (i.e., not instruction tuned) of a high-performing and openly available model, Llama 3 8B (AI@Meta 2024), to fine-tune the model on plain text data. We fine-tune Temporal Adapters with the causal language modeling objective, i.e. predicting next tokens, as shown in Figure 2. We emphasize that the training procedure is completely independent of any survey data and only considers the word tokens that occur in the text data that we obtain from our Twitter panel. In other words, our Temporal Adapters fine-tune the LLM to more closely follow the token probabilities present in the tweets of a given week. We only consider survey data for evaluation, after the training has been completed. Our method aims to fit the LLM to a specific week’s tweets but the evaluation can only be performed across weeks after the training of all weeks has been completed. We thus opt to not have a test set of tweets in each week, as it would not prevent us from over-fitting to tweets from a specific week.
@@
Adapters are a parameter-efficient approach to fine-tuning LLMs that allows researchers to easily swap model properties (Pfeiffer et al. 2023). All types of adapters typically add a very small fraction of additional parameters to a language model with the goal to only train these parameters while keeping the original model weights frozen. This drastically improves training efficiency with little to no loss in training performance, as only relevant parts of the a model can be targeted during training. While many specific implementations of adapters have been proposed, LoRA adapters (Hu et al. 2021) have evolved to be one of the most popular types. They add rank-decomposition matrices to each layer of the transformer architecture. We train LoRA adapters of rank $128$ as a compromise between fine-tuning efficiency and training quality. This is twice the amount of parameters as the authors of the original LoRA paper considered necessary (Hu et al. 2021), but still only $0.67\%$ of the weights that would be trained when fully fine-tuning the model.
@@
We conduct a partial hyperparameter search by training adapters with different learning rates, and for up to 8 epochs. We find that fine-tuning adapters works best for our purposes with a small learning rate of $5*10^{-6}$ and at most $1$ training epoch. The training loss is mostly stable after $0.5$ epochs at an average of $3.4$ across all adapters. In a preliminary experiment with a larger learning rate of $10^{-4}$, training loss decreased to around $2.6$ after $8$ training epochs but the cross-correlations of the attitudes we extracted with survey data were much lower. This is most likely due to the model over-fitting on the training data (tweets) that is quite different from the survey questions we prompt at inference time. We show performance of our method for varying LLM temperatures in Appendix Figure 15, and report results in the main paper for the Llama 3 default temperature if $0.6$. We train adapters with 3 different training seeds for each week to better understand the reliability of our training method. We used bfloat16 and the adamw_torch_fused optimizer with a batch size of $6$ and $4$ gradient accumulation steps. The training was conducted on two NVIDIA H100 GPUs in Distributed-Data-Parallel mode and takes approximately 20 minutes per adapter.
Figure 2: Temporal Adapter Fine-Tuning. We concatenate each week’s tweets into chunks for batch-based fine-tuning. While fine-tuning each week’s parameter-efficient LoRA (Hu et al. 2021) adapter, the original model weights are kept frozen. Fine-Tuning is performed with the causal language modeling task, i.e., next-token prediction on the original tweets, independent of any survey data.
4.2 Extracting Survey Answers
@@
Once we have finished adapter training, we focus on model inference and extract answers to survey questions as follows. We first concatenate a survey question with each of its $n$ answer options, obtaining $n$ separate prompts. We experiment with an optional answer prefix that is build for the particular survey question. For instance, for Britain’s Mood, Measured Weekly, we add the optional answer prefix ‘I felt’ which is followed by the answer options ‘happy’, ‘sad’ etc. This is to accommodate for the fact that first-token probabilities can be unreliable for scoring survey answers from LLMs (Wang et al. 2024b). We prompt each answer option separately and without additional labels as previous research has shown order effects and the tendency for LLMs to prefer certain answer labels (Tjuatja et al. 2024; Dominguez-Olmedo, Hardt, and Mendler-Dünner 2023; Wang et al. 2024b).
@@
We then extract answers from the token probabilities of the LLM as follows. Based on a survey answer scoring method used in previous research (Wu et al. 2024; Naous et al. 2024), we perform a single forward pass of the concatenated question, optional answer prefix, and answer option through the LLM. We gather the token probabilities from the last LLM layer for each token in the answer option after applying softmax. If an answer option consists of multiple tokens, we multiply their probabilities. Our extraction method is deterministic, i.e., if the weights and temperature of an LLM are kept constant, so will be the answer probabilities we extract. Unless otherwise noted, we report results for temperature $0.6$, i.e., the Llama 3 default.
@@
We perform survey answer extraction with each weekly adapter activated separately in the LLM and for each of the $3$ seeds. Since swapping adapters and inference is computationally inexpensive, we extract survey answers after training each adapter for $1$ epoch, as well as after every $50$ steps of training. We obtain a time series for each combination of question, answer options, and set of inference hyperparameters. YouGov presents weekly estimates using a smoothed trend line to reduce random fluctuations due to sampling variability. We follow this approach and apply a $3$ week rolling average. For plotting time series, we apply min-max normalization and show the mean probability for each extracted survey answer across all $3$ seeds.
4.3 Evaluation with Empirical Survey Data
@@
Figure 3: Affect Aggregates Extracted from Temporal Adapters. We extract answer probabilities by prompting a weekly fine-tuned Llama 3 8B with the same question wording as in the survey (YouGov 2024a), and compare them to the respective weekly survey data, as well as to a model that classifies selected emotions in Tweets directly (TweetNLP). The time series are min-max normalized and a $3$ week rolling average is applied. The shaded orange area indicates minimum and maximum LLM answer probabilities across $3$ training seeds. Our results descriptively show in the plot a similar trend of our estimate and the survey data and we find strong positive and significant ($p<0.01$) cross-correlation between LLM probabilites and the survey data. Additional time series are provided in Figures 89, and 10 in the Appendix.
@@
We evaluate our longitudinal macroscopes of emotions and attitudes by cross-correlating them with the respective survey data using Pearson’s correlation coefficient. YouGov publishes weekly survey results on Britain’s Mood, which results in $35$ data points in our observation period that we can use to compute cross-correlation. Results on British attitude towards the NHS and towards Boris Johnson as a Prime Minister are only published monthly, i.e., we are left with 9 relevant data points.
We perform permutation tests with $10.000$ permutations, by shuffling the dates associated with each estimate we obtain from a specific Temporal Adapter, while keeping the survey data intact. This provides us with significance values for our cross-correlations and indicates for which estimates the longitudinal effect matters. In other words, significant results indicate that there are changes in the prevalence of an emotion/attitude over time and that we can successfully track these changes. Constant values or random noise in our estimates and the respective survey data would very unlikely produce significant results.
@@
We compare our results to emotion aggregates from a BERT-based model for emotion detection, TweetNLP (Camacho-Collados et al. 2022). This baseline model will help us differentiate between issues in our training data, e.g. coverage error, or differences in expression of emotion, and measurement error specific to our research setup. TweetNLP was pre-trained on the tweet_eval dataset that consists of tweets labeled with $11$ emotions (Mohammad et al. 2018). The model achieves a macro F1 score of $0.72$ on the tweet_eval test set. We classify each tweet in our weekly training datasets separately and calculate the share of every emotion in every week. Being pre-trained on Twitter data makes TweetNLP particularly suited as a baseline model for our setup. However, only 4 of the emotions classified by TweetNLP match with emotions included in Britain’s Mood, Measured Weekly: ‘fear’/‘scared’, ‘sadness’/‘sad’, ‘joy’/‘happy’, and ‘optimism’/‘optimistic’. This highlights the main issue of previous approaches that use classifiers like TweetNLP: They require expensively labeled training data or hand-crafted dictionaries. Our method, in contrast, does not require pre-defining a construct to be measured, but can be flexibly applied to any survey question. We again cross-correlate the results obtained from TweetNLP with the survey data and perform a permutation test.
4.4 Synthetically Mixed Data
To explore the internal validity of our emotion macroscopes, we perform additional experiments with synthetically mixed data that is labeled as ‘happy’ or ‘sad’. With this setup, we can investigate whether a different prevalence of emotion signals indeed corresponds to different affect aggregates obtained from our method. We hypothesize a linear relationship between the amount of ‘sad’ tweets in the training data and the token probability for ‘sad’ as an answer option, and vice versa for ‘happy’. There are $1163$ tweets labeled ‘happy’ and $1326$ tweets labeled ‘sad’ in the tweet_eval dataset. From these tweets, we select $1163$ using $11$ splits: $100\%$ happy + $0\%$ sad, $90\%$ happy + $10\%$ sad, …, and $0\%$ happy + $100\%$ sad. We then train adapters on each of these splits using the same hyperparameters as described above. We repeat the splitting and training procedure using $10$ random seeds to improve robustness. Since the synthetically mixed training datasets are much smaller than our weekly Twitter datasets, the number of training steps/epochs is not directly comparable.
5 Results
@@
Figure 3 shows time series for two collective emotions, scared and happy, that we extract using weekly Temporal Adapters. We use the original survey question developed by YouGov (2024a) and compare our results with nationally representative survey data. We cap the training steps such that we have the same amount of training data across weeks. The plot shows the mean answer probability across $3$ training seeds, with minimum and maximum probabilities, after a $3$ week rolling average was applied and after the time series were min-max normalized. Time series for other emotions are provided in Figures 8 and 9 in the Appendix.
@@
Overall, we observe similar trends between our estimates of collective emotion and the survey data. Scared spikes on the March 23rd, 2020, when the first UK lockdown in response to the COVID-19 pandemic was announced. [^2] https://www.theguardian.com/world/2020/mar/23/boris-johnson-orders-uk-lockdown-to-be-enforced-by-police, Accessed Sept 8th, 2024 Another increase can be seen in our estimate of collective fear around June 1st 2020, at a time when lockdown restrictions were partially lifted.333for England, see for instance https://www.legislation.gov.uk/uksi/2020/558/made, Accessed Sept 8th, 2024 Fear also sees an increase around Christmas 2019. The survey data similarly indicates the largest level of collective fear around the announcement of the lockdown and a small increase around Christmas, but the second spike around June 1st, 2020 is not observed in the survey. This is likely due to sampling errors in the social media data, i.e., more concerned people being active on Twitter, or due to differences in the expression of emotions on social media and in surveys. TweetNLP shows the same second spike in 2020 when used to classify the same training data, as shown in Figure 3.
@@
For collective happiness, we similarly extract estimates with good face validity. Happiness decreases rapidly in the two weeks before the nationwide lockdown was announced and recovers after another two weeks, presumably when people acclimatized with the new situation. This follows the trend that we observe in survey data. Similar to what we have observed for ‘scared’, our estimate deviates from survey data around June 1st 2020, when restrictions were partially lifted. TweetNLP extracts the lowest happiness after June 1st. Our method, instead, more closely tracks the YouGov survey data and identifies the lowest happiness around March 23rd. It should be addressed that we observe larger confidence intervals across $3$ training seeds, and larger week-by-week fluctuation for ‘happy’ as compared to ‘scared’. One possible explanation for this would be that there is less information about happiness in our training data, which would lead to less influence on the fine-tuned adapters, and thus result in larger confidence intervals after min-max scaling. Another explanation could be that information about happiness is less evenly distributed in our training data. We sample an equal amount of text for each week and randomly shuffle the training data, so a skewed distribution of information about happiness would increase the relevance of different training seeds. Thirdly, ‘happy’ could be a more difficult to learn concept for LLMs, potentially because it is expressed with more diverse language than ‘scared’, leading to more error surrounding the concept in our outcome estimate.
@@
Figure 4: Several Extracted Emotions Highly Correlate with Survey Data. We cross-correlate the answers we extract from Llama 3 with the respective British survey data (YouGov 2024a). Across $3$ training seeds, we show minimum and maximum correlation with error bars and indicate the worst $p$ value (*$p<0.05$,**$p<0.01$,***$p<0.001$). Our results vary strongly between emotions, but they are in line with TweetNLP’s estimates. As opposed to TweetNLP, our method can be flexibly applied to extract any emotion as it does not require labeled training data.
We calculate cross-correlations (Pearson) between the longitudinal data we extract and the respective survey data, and perform a significance test with $10,000$ permutations. Figure 4 shows the results for all emotions found in Britain’s Mood, Measured Weekly across 3 training seeds, as well as the cross-correlation for our baseline method, TweetNLP (Camacho-Collados et al. 2022). We observe three groups of emotions: (i) emotions that show a strong positive, significant correlation, (ii) emotions that are weakly positively or negatively correlated, and (iii) optimistic, which is strongly negatively correlated. In the first group, we find both positively and negatively connoted emotions, and in particular affective phenomena that are sometimes referred to as being part of a set of basic emotions: ‘happiness’, ‘fear’, and ‘sadness’ (Ekman 1992). Significance levels obtained from a permutation test indicates that there are non-random changes in the emotions over time and that we can successfully track these changes. For collective emotions where an estimate can be extracted from TweetNLP, i.e., where labeled training data was available, we find that our method produces comparable, yet slightly worse results. We did not expect to outperform TweetNLP as it classifies individual tweets, while the Temporal Adapters in our setup have to aggregate information from many tweets first, before survey answers are predicted. We do, however, obtain strong positive, significant correlations for emotions for which no TweetNLP estimate is available: ‘frustrated’, ‘bored’, and ‘energetic’. This flexibility of our approach is a major benefit over existing methods such as TweetNLP and is further highlighted in the additional results reported below. For group (ii), emotions with weak positive/negative correlations, it remains unclear why these collective emotions are hard to estimate from the data that we have available. One possible explanation might be that in contrast to the affective phenomena found in group (i), group (ii) contains more phenomena that clearly fall under ‘mood’ rather than ‘emotion’, i.e., they are more diffuse and longer lived, which might contradict typical social media behavior. Finally, we find a significant strong negative cross-correlation for ‘optimistic’, both for our estimate as well as for the baseline model. This points to general differences in the expression of this emotions between surveys and social media.
mean SE median SE SE std. dev.
frustrated 0.0414 0.0185 0.0507
scared 0.0509 0.0108 0.0872
bored 0.0387 0.0141 0.0586
sad 0.0502 0.0186 0.0666
energetic 0.0819 0.0531 0.0848
happy 0.0780 0.0482 0.0763
lonely 0.1229 0.0908 0.1256
inspired 0.1365 0.0737 0.1628
stressed 0.2284 0.0905 0.2865
content 0.1899 0.0806 0.2179
don’t know 0.1637 0.1439 0.1365
apathetic 0.2171 0.0987 0.2286
other 0.3119 0.3021 0.2347
optimistic 0.3595 0.3598 0.2565
Table 1: Means and Standard Deviations in Squared Errors (SE). We individually min-max-normalize each time series – survey data and Llama 3 estimates from 3 training seeds – before we calculate squared errors between the same points in time in the YouGov survey data and the Llama 3 estimates. We find that strong positive cross-correlation shown in Figure 4 coincides with low mean squared error.
For some collective emotions, we observe variable performance across time, i.e., our estimates obtained from Temporal Adapters closely track the YouGov survey data only for a certain time period. To address this issue, we additionally calculate squared errors between our estimates and the YouGov survey data, and report summary statistics of these squared errors in Table 1. Overall, we observe that strong positive cross-correlation for collective emotions such as frustrated, bored, or scared coincides with low mean squared errors and low standard deviations of these squared errors.
@@
We create synthetically mixed training data from the tweet_eval dataset (Mohammad et al. 2018) in 11 splits ranging from $100\%$ sad to $100\%$ happy. We then prompt the model with the same question as used in Britain’s Mood, Measured Weekly and extract answer probabilities for the answer options ‘happy’ and ‘sad’. We hypothesize a positive linear correlation between the share of ‘happy’ tweets in the training data and the probability for answer option ‘happy’, and vice versa for ‘sad’. Figure 5 shows the mean and standard deviation of extracted answers across 10 training seeds. We find strong positive correlations for both answer options, which supports the internal validity of our method – more ‘happy’ tweets actually lead to a higher answer probability for ‘happy’ and vice versa for ‘sad’. For both emotions, the relationship between training data and extracted answer is surprisingly linear. Our answer scoring method does not ensure that the probabilities across all answer options add up to $1$, since we apply softmax on the last layer of the LLM, and each answer probability depends on probabilities of all other tokens in the vocabulary. The error bands indicate substantial random error resulting from the training and underline the importance of evaluating results obtained from multiple seeds.
@@
Figure 5: Internal Validity Demonstrated in Experiments with Synthetically Mixed Data. We synthetically mix LLM training data with splits ranging from data that is labeled $100\%$ sad to $100\%$ happy. We then extract answers to YouGov (2024a)’s survey question at each split, and show mean and standard deviation over 10 training seeds. The results support the internal validity of our extraction method, by showing a surprisingly linear relationship between training data ratio and extracted estimate, but also random training error.
@@
In preliminary experiments, we investigated different hyperparameters for model training and for the extraction of affect aggregates, see Figure 16 in the Appendix. We investigated cross-correlations after different numbers of training steps and epochs and found little benefit in fine-tuning the model for more than $1$ epoch. While longer fine-tuning further improves training loss in the causal language modeling objective, the highest cross-correlations with survey data are observed relatively early on. This is likely due to the language model overfitting on the training data, which itself is quite different from the questions that we use for answer extraction. Similarly for learning rates, we found that smaller learning rates generally perform better, as long as loss does not remain constant. We find $5*10^{-6}$ to work best on our data. Finally, we found that training all adapters with the same amount of training data generally works better than training all adapters for $1$ epoch on all the data that was available in this week. In other words, we extracted answers for all Figures presented previously in this section after training each adapter for $350$ steps instead of $1$ epoch.
For answer extraction, we investigated the effect of temperature and of whether or not an answer prefix was attached to the answer options. A lower temperature leads to less evenly distributed token probabilities, similar to ‘enhancing contrast’. We found that this is beneficial when extracted attitudes have low noise, but that it can also increase random fluctuation. We also found that lower temperatures tend to overemphasize some answer options. Finally, adding an answer prefix leads to more consistent results. This is in line with previous findings on first-token probabilities (Wang et al. 2024b) and is likely related to LLMs being trained to create text rather than to answer surveys with only an answer option.